40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 121 of 330  Not logged in ELOG logo
ID Date Author Type Category Subject
  10531   Wed Sep 24 11:02:38 2014 manasaUpdateLSCMoving SRM

I looked at the CAD layout and it seems like we will clearly be clipping POY if we move SRM by 7.5cm. Since POY is not visible at low power, we cannot be sure about the clipping.

We should have a plan B before we move everything. I suggest we move a combination of SRM and SR2 to get the desired SRC length.
Moving SR2 will require extra effort to walk the beam unclipped through all the 6 output steering mirrors that follow; but there will be little room for error if we use irides to propagate the beam through the first 4 mirrors that are in the BS and ITMY chamber.

  10530   Wed Sep 24 08:40:29 2014 SteveUpdateVACVent has started

 

 Jam nuts checked. Oplev servos turned off. Particle count checked. Vertex crane functionality checked.

Attachment 1: vent77.png
vent77.png
Attachment 2: bvent77.png
bvent77.png
  10529   Wed Sep 24 08:39:32 2014 Jenne, EricQUpdateGeneralVent prep for SRC length change

Quote:

As Q mentioned in elog 10527, (prompted by Koji's email this afternoon) we are prepping the IFO for vent.  Here is a copy of the pre-vent checklist from the wiki, updated as we work:

 

Pre-vent checklists

 
  1. Center all oplevs/IPPOS/IPANG
  2. Align the arm cavities for IR and align the green lasers to the arms.
  3. Make a record of the MC pointing
  4. Align the beam at the PSL angle and position QPDs
  5. Reduce input power by adjusting wave plate+PBS setup on the PSL table BEFORE the PMC. (Using the WP + PBS that already exist after the laser.)
  6. Replace 10% BS before MC REFL PD with Y1 mirror and lock MC at low power.
  7. Close shutter of PSL-IR and green shutters at the ends
  8. Make sure the jam nuts are protecting bellows

Notes:

1 & 2:  Locked arms on IR, ran ASS.  Unlocked IFO, aligned PRM for good POP flashes, aligned SRM for symmetric AS flashes.  Aligned all oplevs.  Used PZTs to align Xgreen to arm. Used knobs to align Ygreen to arm.  With PS:L green shutter closed, Xgreen  = 0.520, Ygreen = 0.680.

3:  Moved MC servo output cable that goes to ADC from OUT2 (which we had been using for monitoring AO path signals) back to its usual OUT1 (which is MC_L).  This is used in the spot position measurement script.  Spots at:  [2.32, -0.50, 1.97, -1.11, 0.26, -1.86] mm.

4: Done -Q

5:  Removed a PD that was monitoring the light coming backwards through the Faraday that sits just after the laser, just in case (confirmed that beam dump behind PD was catching beam).  Other port of PBS just had regular black hole dump.  Adjusted half wave plate until we had ~90mW just before injection into the vacuum.

6: Completed. Locked MC manually at transmission of ~1150, but low power autolocker isn't working. This isn't a critical thing, and can be fixed at any point during the vent. -Q

7: Shutters closed. Ready for Steve to check nuts and begin venting! -Q

 

  10528   Tue Sep 23 17:56:13 2014 Jenne, EricQUpdateGeneralVent prep for SRC length change

As Q mentioned in elog 10527, (prompted by Koji's email this afternoon) we are prepping the IFO for vent.  Here is a copy of the pre-vent checklist from the wiki, updated as we work:

 

Pre-vent checklists

 
  1. Center all oplevs/IPPOS/IPANG
  2. Align the arm cavities for IR and align the green lasers to the arms.
  3. Make a record of the MC pointing
  4. Align the beam at the PSL angle and position QPDs
  5. Reduce input power by adjusting wave plate+PBS setup on the PSL table BEFORE the PMC. (Using the WP + PBS that already exist after the laser.)
  6. Replace 10% BS before MC REFL PD with Y1 mirror and lock MC at low power.
  7. Close shutter of PSL-IR and green shutters at the ends
  8. Make sure the jam nuts are protecting bellows

Notes:

1 & 2:  Locked arms on IR, ran ASS.  Unlocked IFO, aligned PRM for good POP flashes, aligned SRM for symmetric AS flashes.  Aligned all oplevs.  Used PZTs to align Xgreen to arm. Used knobs to align Ygreen to arm.  With PS:L green shutter closed, Xgreen  = 0.520, Ygreen = 0.680.

3:  Moved MC servo output cable that goes to ADC from OUT2 (which we had been using for monitoring AO path signals) back to its usual OUT1 (which is MC_L).  This is used in the spot position measurement script.  Spots at:  [2.32, -0.50, 1.97, -1.11, 0.26, -1.86] mm.

4: Done -Q

5:  Removed a PD that was monitoring the light coming backwards through the Faraday that sits just after the laser, just in case (confirmed that beam dump behind PD was catching beam).  Other port of PBS just had regular black hole dump.  Adjusted half wave plate until we had ~90mW just before injection into the vacuum.

6: Completed. Locked MC manually at transmission of ~1150, but low power autolocker isn't working. This isn't a critical thing, and can be fixed at any point during the vent. -Q

7: Shutters closed. Ready for Steve to check nuts and begin venting! -Q

  10527   Tue Sep 23 17:37:10 2014 ericqUpdateLSCDRMI locking

Rather than using a CAD drawing, I used Gabriele's code from ELOG 9590 to try and judge if we could shorten the SRC by the appropriate length, without clipping the SR3-SR2 beam. 

Specifically, I used these lines:

% Move SRM 7.5 towards SR2, parallel to beam

delta=75;

dAS = BS2-AS; Vector from SRM to SR2

dASmag = sqrt(dAS(1)^2+dAS(2)^2);

dMove = delta*dAS/dASmag;  delta times unit vector

CS = CS+dMove;

draw_sos(CS, 180/pi*angles)

 
to help generate this plot:

SRCcorrection.pdf

 


As a reminder, Gabriele's code used the following logic:

  • We know the nominal dimensions of all of the suspensions
  • We hand measured various distances between features of the suspension structures. (Corner to corner)
  • A global fit, minimizing the maximum error, reconstructed the positions of the suspensions. 
  • Beam positions assumed to be ideally aligned. 
  • Beam trajectories traced out, and optical path lengths estimated (taking into account changing indices of refraction due to flipped mirrors)

In my opinion, this is the best estimate of beam trajectory that we currently have.

Thus, from looking at the plot above, I claim we can correct the SRC length without clipping the beam by moving the SRM forward by the required 7.5cm.

Although the measured distance may be off on the order of a cm (since our PRC correction had a 0.5cm disagreement between interferometric and hand distance measurements), this will nevertheless markedly improve our 3F DRMI sensing, based on my previous ELOG. 

 


Hence, given our discussions last week, Jenne and I will proceed to ready the interferometer for venting in the morning, by following the vent checklist.

Our sole objective for this vent is this move of the SRM. 

Steve, please check the jam nuts, and begin the vent when you get in.  Thanks!

  10526   Tue Sep 23 09:37:12 2014 SteveUpdateSUSIOO and temp changes
Attachment 1: IOO.png
IOO.png
  10525   Tue Sep 23 08:34:14 2014 SteveUpdateSUSHeNe laser test

 

 The room temp drops 1 degree C on the 4th day. The weather has changed.

Attachment 1: tempDrop.png
tempDrop.png
  10524   Mon Sep 22 15:20:32 2014 SteveUpdateVACRGA scan at day 229

 

 

Attachment 1: RGA@229day.png
RGA@229day.png
Attachment 2: RGAscan229d.png
RGAscan229d.png
  10523   Mon Sep 22 10:18:58 2014 steveUpdatePEMlab temperatures

 

 

Attachment 1: summerheat.png
summerheat.png
  10522   Mon Sep 22 10:14:07 2014 SteveUpdateSUSHeNe laser test

Quote:

TEST QPD sn 222 was calibrated with 1103P directly looking into it from 1 m. ND2 filter was on the qpd.

 

 

Attachment 1: 1103Pmon3d.png
1103Pmon3d.png
  10521   Fri Sep 19 13:12:07 2014 JenneUpdateLSCAO path glitches

Quote:

Discontinuities / glitches could be seen in the CM board fast output when MC board gains were changed, which isn't so nice. Incidentally, I notice now that each lock loss corresponded to a step of AO gain on the CM board.

Back in May I looked at all the glitches that happen when we change the AO gain slider on the CM board - see elog 9938.   I wonder if the MC IN2 gain slider has the same issues.  I think I'll look at this this afternoon. Maybe we can set the CM board gain someplace, and just use the MC IN2 slider (if it's not as glitchy) for the delicate part where we're just about to cross unity, and then later we can again use the CM board's AO gain.

EDIT:  Yes, the glitches on the CM board AO path are *much* bigger, and more frequent.  Interestingly, the biggest glitches were every 4 dB.  When I went from -29 to -28, again from -25 to -24, -21 to -20, etc.  I saw the largest glitches on the MC IN2 slider going -29 to -28 and -17 to -16, but if there were small glitches at other transitions, they didn't hit my trigger levels.  I think next time I try engaging the AO path I'll try to do the delicate stuff by upping the MC IN2 gain rather than the CM board AO gain.

  10520   Fri Sep 19 04:05:05 2014 ericqUpdateLSCAO path partly engaged

More AO efforts. No huge news. 

Came at AO from each side. For each sign, I lost lock just a few dB from the AO portion of the loop crossing unity gain. Both attempts were about arm powers of 1, which should correspond to ~300pm CARM offset, which I have simulated the crossover as possible with my current loop models (including latest MC loop). The gain steps were usually 6dB in between measurements. 

Positive polarity on CM board screen:

I made it to +5 dB of the last plot here, but the 6th broke it open. Gains on CM In2, CM AO, and MC In2 were -6, -4, -2 on that last, lock breaking, step. 

TFSR785_19-09-2014_020555.pdf

Negative polarity on CM board screen:

Lost it just 2dB above the last trace. Gains were -6, +1, -2 (So, overall 5dB higher than the other polarization)

TFSR785_19-09-2014_033920.pdf

Many things happened in between these two lock stretches, but I'm not sure what may or may not have affected things. They include:

  • Jenne mentioned PRMI being fussy earlier in the evening. I adjusted REFL33 and POP22 angles during a PRMI lock, while CARM was held away with ALS. My simulations suggest that there are small changes to the 3F sensing when the arms are totally absent, but doing it at a finite CARM offset is closer to where we want it, it seems. 
  • I tried using REFL165Q for MICH, since my simulations suggest a better MICH/PRCL angle, which would stave off cross couplings. Lined up excitations, etc., but no luck. 
  • I measured the PRMI loops
    • found PRCL to have ~200Hz UGF, 8dB gain peaking. Maybe a little high, but didn't seem terrible. 
    • MICH had UGF of around 20Hz, with the FM gain at 0.8. By the shape of the phase bubble, the loop seems designed for higher bandwidth. I raised the gain to 2.5 for a 70ishHz UGF, and called in FMs 7 and 9 for additional triggered boosts. Things seemed to stay locked pretty well. 
  • Lower excitation amplitude the second time around, measuring the AO loop. Looking at the CM output spectra, you can see the excitation wailing away; I wanted to avoid it.

The location of the CARM resonance peak lines up with my simulation, which is good, but there appears to be less phase than expected... I tried making sure that we don't have any whitening uncompensated for, but it looked ok. All my AO path loop model contains is the CM board TF (measured and fitted), the IMC seen as an actuator(measured and fitted), and the REFLDC optical TF (simulated in MIST). Maybe the DC path of whatever diode this is coming from needs to be included...

Discontinuities / glitches could be seen in the CM board fast output when MC board gains were changed, which isn't so nice. Incidentally, I notice now that each lock loss corresponded to a step of AO gain on the CM board.

  10519   Thu Sep 18 17:44:55 2014 JenneUpdateLSCOld AO cable pulled

[Q, Jenne]

We pulled the old 2-pin lemo cable after I had a look at the connectors.  When I unscrewed the connector on the MC side, one of the wires came off.  I suspect that it was still hanging on a bit, but my torquing it finally killed it. 

We pulled the cable with the idea of resoldering the connectors, but there are at least 2 places where the cable has been squished enough that the shielding or the inner wires are exposed.  These places aren't near enough the ends to just cut the cable short.

Downs doesn't have a spool of shielded twisted single-pair cable, so Todd is going to get me the part number for the cable they use, and I've asked Steve to order it tomorrow. 

For now, we will continue using the BNC cable that we installed last night - I don't think it's worth resoldering and putting in a crappy 2-pin lemo cable that we'll just throw out in a week.

  10518   Thu Sep 18 10:08:07 2014 steveSummarySUSoplev laser summary updated

 

Quote:

 

Quote:

 

                  2005              ALL oplev servos use Coherent DIODE LASERS # 31-0425-000, 670 nm, 1 mW

    Sep. 28, 2006              optical lever noise budget with DC readout in 40m,  LIGO- T060234-00-R, Reinecke & Rana

    May  22, 2007              BS, SRM & PRM  He Ne 1103P takes over from diode

    May  29, 2007              low RIN He Ne JDSU 1103P selected, 5 purchased sn: T8078254, T8078256, T8078257, T8078258 & T8077178 in Sep. 2007

    Nov  30, 2007               Uniphase 1103P divergence measured

    Nov. 30, 2007               ETMX old Uniphase 1103P  from 2002 dies: .............., running time not known......~3-5 years?

    May 19, 2008               ETMY old Uniphase 1103P from 1999 dies;.....................running time not known.....~    ?

    Oct.  2, 2008                ITMX & ITMY are still diodes, meaning others are converted to 1103P earlier

 

                     JDSU 1103P were replaced as follows:

   May 11, 2011                ETMX replaced, life time 1,258 days  or 3.4 years

   May 13, 2014               ETMX , LT 1,098 days or 3 y

   May 22, 2012               ETMY,  LT 1,464 days or  4 y

   Oct.  5, 2011                BS & PRM, LT 4 years,  laser in place at 1,037 days or 2.8 y

   Sep. 13, 2011               ITMY  old 1103P &    SRM    diode laser replaced by 1125P  ..........old He life time is not known, 1125P in place 1,059 days or 2.9 y

   June 26, 2013              ITMX 622 days or 1.7 y    note: we changed because of beam quality.........................laser in place 420 days or 1.2 y

 

  Sep. 27, 2013               purchased 3 JDSU 1103P lasers, sn: P893516, P893518, P893519 ......2 spares ( also 2 spares of 1125P of 5 mW & larger body )

 

      May  13, 2014             ETMX,  .............laser in place 90 d

      May  22, 2012             ETMY, 

     Oct.  7,  2013             ETMY,  LT  503 d  or  1.4 y............bad beam quality ?

     Aug. 8,  2014              ETMY,  .............laser in place   425 days  or  1.2 y

 

      Sept. 5, 2014              new 1103P, sn P893516  installed at SP table for aLIGO oplev use qualification

     

  10517   Thu Sep 18 09:17:47 2014 SteveUpdateSUSHeNe laser test

TEST QPD sn 222 was calibrated with 1103P directly looking into it from 1 m. ND2 filter was on the qpd.

 

Attachment 1: calibrated_16hrs.png
calibrated_16hrs.png
Attachment 2: QPD222cal.jpg
QPD222cal.jpg
Attachment 3: 1103P40s40min.png
1103P40s40min.png
  10516   Thu Sep 18 02:42:28 2014 JenneUpdateLSCAO path partly engaged

Tonight was a night of trying to engage the AO path.  The idea was to sit at arm powers of a few on sqrtInvTrans for CARM and ALS for DARM, and try to increase the gain for REFLDC->AO path.

No exciting nit-picky details in locking procedure.  Mostly it was just a night of trying many times. 

The biggest thing that Q and I found tonight was that the 2-pin lemo cable connecting the CM board's SERVO OUT to the MC board's IN2 is shitty.  The symptom that led to this investigation was that I could increase the AO path gain arbitrarily, and have no change in the measured analog CM loop transfer function. We checked that the CM board servo out spit out signals that were roughly what we expected based on our ~2kHz excitation.  However, if we look at digitized signals from the MC board, the noise level was very high, with loads of 60Hz lines, and a teensy-tiny signal peak.  We put a small drive directly into the MC board and could see that, so we determined that the cable is bad.  We have unplugged the white 2-pin lemo, and ran a long BNC cable between the 2 boards.  Tomorrow we need to make a new 2-pin lemo cable so that we can have the lower noise differential drive signal.

After putting in the temporary cable, we do see an excitation sent to the CM board showing up after the MC board.  For this monitoring, the MC_L cable to the ADC has been borrowed, so instead of being the OUT1, the regular length signal, MC_L is currently the OUT2 monitor right after the board inputs. 

At some point in the evening, around 1:15am, ETMX started exhibiting the annoying behavior of wandering off sometimes.  I went in and pushed on the SUS cables to the satellite box, and I think it has helped, although I still saw the drift at least once after the cable-squishing.

Other than that, it has just been many trials.

The best was one where I was holding the arm powers around 4, and got the CM board's AO gain to -8 dB and the MC board's IN2 AO gain to -4 dB. I lost lock trying to increase the CM board gain to -7 dB. 

I took several transfer functions, and used Q's nifty "SRmeasure" script to gather data, and Q made a plot to see the progress.

TF progress plots:

0020.pdf

Time series of that lockloss:

Zoom_TRXTRY4_EngagingAO_1095059162.png

I don't know yet if the polarity of the CM board should be plus or minus.  This series was taken with "minus".  But,  since the phase looked opposite of Q's single arm CM board checkout from several months ago, we did a few trials with the polarity switched to "plus".  I thought we weren't getting as high of AO path gains, so I switched back to "minus", but the last few trials didn't get even as far as the plus trials did.  So, I still don't know which sign we want.

  10515   Wed Sep 17 18:36:03 2014 KojiHowToGeneralHow to run DTT measurement automatically
  • Suppose you have a dtt template name test.xml
  • The file test.dtt

    open
    restore test.xml
    run -w
    save test2.xml
    quit
     
  • Run diag < test.dtt
  • The result is saved in test2.xml
  10514   Wed Sep 17 15:40:00 2014 ericqUpdateLSCDRMI locking

I have not had any success the past two days in getting an interferometric measurement of the SRC length. 

So, the question posed at today's meeting was: "How precisely do we need to change the SRC length to be able to lock the DRMI on 3F"

The two ways I could think to quantify this are:

  • How much MICH -> [S,P]RCL cross coupling is ok?
  • How much [S,P]RCL ->  MICH cross coupling is ok?

REFL33 should have its phase set to put PRCL along I, and REFL165 should have SRCL along I, so the simulation result that matters is the angle of MICH in these planes. The cross couplings are then given by the appropriate trigonometric projections. In the following plots, I used 10% as the acceptable cross coupling in either direction. 

Result:

dcAngleDiff_srcL.pdf

Thus,

  • To limit the MICH -> [S,P]RCL coupling to 10%, we must hit the ideal length within +- 1.2cm.
  • To limit the SRCL -> MICH coupling to 10%, we must hit the ideal length within +- 2mm.
  • It doesn't look like we can get the REFL33 angle totally to 90 degrees, REFL165 looks more promising.

Code (finesse + pykat + ipython notebook) and plots are attached. 

Attachment 2: drfpmiVertexSensing.zip
  10513   Wed Sep 17 13:41:00 2014 JenneUpdateSUSNew calibrated channels for test QPD

Steve asked about calibrating the QPD, so I set up some new epics records so that we can have calibrated versions of the QPD output.

The new channels are called C1:ASC-TESTQPD_Y_Calc and C1:ASC-TESTQPD_X_Calc for pitch and yaw, respectively.


Details:

 * I modified /cvs/cds/caltech/target/c1iscaux/QPD.db to add 2 new channels.  Since we are currently plugged into the IPPOS channels, I didn't want to modify the units of IPPOS, which is why I created new channels.  The new channels are just the IPPOS normalized X and Y channels, multiplied by a calibration factor.  Steve has already done a rough calibration for his setup, so I used those numbers (0.15 urad/ct for pitch and 0.25 urad/ct for yaw). 

* Rebooted c1iscaux.  This required adding it to chiara's /etc/hosts file.

* Added the channels to the /opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini file so that the channels would appear in dataviewer.

* Restarted the framebuilder daqd process.


How to modify the calibration:

1) On a control room workstation, cd /cvs/cds/caltech/target/c1iscaux to get to the right folder.  (Note that this is still in the old cvs/cds place, *not* the new opt/rtcds place)

2)  open the epics database file by typing sudo emacs QPD.db.  Since this is a protected file, you need to use the "sudo" command, and will have to type in the usual controls password. 

3)  Find the "records" that have the channel names C1:ASC-TESTQPD_Y_Calc and C1:ASC-TESTQPD_X_Calc by scrolling down.  (Right now they are on lines #550 and #561 of the text file).

4) For each of these 2 records, modify the calibration in the line that says something like field(CALC,"(A*0.25)").  In this example, the current calibration is 0.25 urad/oldCount.  Change the number to the new value.

5) Save the file.  If you followed the procedure in step2 and used the emacs program and you can't use the mouse, do the following:  Hold down the "ctrl" key.  Keeping ctrl pushed down, push the "x" key.  Still keeping ctrl pushed down, push the "s" key. 

6) Close the file.  If you followed the procedure in step2 and used the emacs program and you can't use the mouse, do the following:  Hold down the "ctrl" key.  Keeping ctrl pushed down, push the "x" key.  Still keeping ctrl pushed down, push the "c" key. 

7) Reboot the slow computer called c1iscaux.  You should be able to do this remotely by typing telnet c1iscaux, and then typing reboot.  If that doesn't work, you may have to go into the IFO room and power cycle the crate by turning the key.  This computer is in 1Y3, near the bottom.

8) Check that you can see your channels - you should be finished now!

For steps 3 and 4, here is a screenshot of the lines in the text file:

NewQPDcalibratedChans.png

  10512   Wed Sep 17 01:40:55 2014 JenneUpdateLSCCloser to REFL DC? Maybe?

I tried a bunch of times to reduce my CARM offset so I could jump to REFLDC digitally, but I think I'm maybe being a little ambitious with the arm power I'm trying to get to.

I have modified the carm_cm_up script so that it does my new procedure.  Everything is the same through locking the PRCL and MICH on 3f.  Then it reduces the CARM offset to 1.5 nm.  This is where we *used* to transition to sqrtInvTrans.  Now I have it going a bit farther to 0.5 nm, and arm powers of about 1 before doing that transition.  Also, before it transitions it lowers the CARM gain and engages the 1kHz lowpass in FM9.  A gain of about 4 is fine to keep the gain peaking in the CARM loop to only about 10dB, and sets a UGF of 100Hz which is the peak of the phase bubble with the lowpass engaged. 

Once I got to this point (several times tonight), I turned on CARM and DARM oscillations and looked at the transfer functions between (CARM and REFLDC) and (DARM and AS55Q). I have 2 DTT templates setup for this, in /users/Templates/PRFPMI.  These templates assume that you have your new DARM signal (AS55) going to SRCL_IN1 and your new CARM signal (REFLDC, which is actually REFL11I coming through the CM board) going to MC_IN1. 

I'm not sure why I'm losing lock. I don't see anything terribly telling on the time series plots, in particular none of the loops look like they are oscillating.  Here is one of the better examples from this evening:

Zoom_TRXTRYupto25_1094967592.png

Other notes:

* I realigned the Xgreen on the PSL table (again) to maximize the beatnote amplitude.  Y was fine, but X was very poorly overlapped on the camera.

* I put the SR785 back by the LSC rack and plugged it into the CM board for transfer functions.  Didn't take any tonight.

* We have a small wishlist for scripting things:  (1) DRMI restore script should reset REFL11 to "normal" REFL11.  (2) CARM/DARM acquisition restore script should reset REFL11 to REFLDC.  (3) CARM/DARM acquisition restore should also set PRMI parameters (as Q noted last week).

  10511   Tue Sep 16 17:02:41 2014 SteveUpdateComputer Scripts / ProgramsPSL output shutter is floating

Quote:

Quote:

Q and Steve will follow elog 10028 entry to prepare the vacuum system for safe reboot

Here's the sequence of the morning so far:

  • I aligned the IFO (IR arms with ASS, X green with PZTs, PRM with PRMI locked on REFL33)
  • I closed the PSL shutter, and went inside to align PRM and both ITM oplevs (all others were within 10urad of zero in both directions)
  • While aligning those oplevs, I noticed the smell of burnt electronics. We tracked it down to the +15V sorensen in the rack nearest the PSL table
    • I claim the precipitating event was PSL shutter activity. If I recall correctly, the seismic rainbow traces went bonkers around the same time as the shutter was closed. There is a Guralp interface in the rack powered by the failed sorensen, so this would explain the erratic seismometer signals correlated with the power supply failure. We will look into potential shorts caused by the shutter. (Steve looked up the PMC trans and Guralp DQ channels, and confirmed the temporal coincidence of the events.)
  • We shut off all of the sorensens so that electronics were not being driven asymmetrically. 
  • Steve and I secured the vacuum system for computer reboots, as referred to in Steve's elog. Some combination of Jenne, Rana and Manasa shut down the control room computers, and turned off the watchdogs. 
  • Manasa and I moved Chiara inside, next to Mafalda, along with its backup HDs. It has been labeled. 
  • Booted up control room machines, they came up happy. 
  • FB and front-ends didn't need reboot, for some lucky reason. Watchdogs came back happily, oplev spots didn't move noticeably. 

The IFO is still down, as the PMC won't lock without the rack power, and we haven't pinned down the shorting mechanism. We don't want the replacement sorensen to immediately blow when plugged in. 

 The Uniblitz shutter is insulated at the optical table. I only realized it when it was in my hand. The 3"x2" base plate is black DELRIN !

 In the future we must label black Delrin base plate where it is used. Now we have  white Delrin and light bran PEEK  base plates  for the same function.

  10510   Tue Sep 16 16:03:36 2014 KojiUpdatePSLLaser turned on

Quote:

 Our janitor turned off the laser accidentally. 

 Didn't you take this opportunity to replace the cooling fan of the innolight controller?

  10509   Tue Sep 16 14:26:45 2014 ericqUpdatePSLLaser turned on

Quote:

 Our janitor turned off the laser accidentally. 

 The PMC wasn't locking very happily after this. I tweaked the pointing onto the PMC REFL diode, to make sure it was centered, and touched the alignment into the PMC. I also reset the FSS Slow output to zero. It took a little while for the laser to settle in, for some reason, but the transmission is up at 0.80 now. 

Tweaked MC2 pointing to get the MC transmission high enough to let WFS kick in, which nicely got the rest of the MC alignment done. After that, I offloaded the WFS into the MC suspensions. 

Lastly, I ran the command that Rana posted in ELOG 10391, to set the FSS input offset (From -0.18 to -0.06)

  10508   Tue Sep 16 10:47:52 2014 SteveUpdatePSLLaser turned on

 Our janitor turned off the laser accidentally. 

  10507   Mon Sep 15 18:55:51 2014 ranaUpdateDAQ40m frames onto the cluster

 Dan Kozak is rsync transferring /frames from NODUS over to the LDAS grid. He's doing this without a BW limit, but even so its going to take a couple weeks. If nodus seems pokey or the net connection to the outside world is too tight, then please let me and him know so that he can throttle the pipe a little.

  10506   Mon Sep 15 15:52:44 2014 SteveUpdateSUSPRM damping restored

 The PRM side was kicked up

  10505   Mon Sep 15 14:37:49 2014 SteveUpdateGeneralOphir pmeter has no filter already

Quote:

Ophir power meter gets new filter with calibration. This is not cheap. It was the second time we lost it.

Filter leash is attached.

 Some one already took  off the filter and did not care to put it back on. This is carelessness!

  10504   Fri Sep 12 16:30:48 2014 SteveUpdateSUSHeNe laser test preparation

Quote:

IP POS cable was swapped with old SP-QPD sn222 at the LSC rack.  So there is NO IP POS temporarily.

This QPDsn222 will be used the HeNe oplev test for aLIGO

 

 QPDsn222 is on translation stage with ND2 filter on SP table. The 1103P is mounted with two large V mounts 1 m away.

This qpd will be calibrated Monday. It has only slow outputs.

Attachment 1: ND2.png
ND2.png
Attachment 2: qpd222ND21103P.jpg
qpd222ND21103P.jpg
  10503   Fri Sep 12 15:10:09 2014 JenneUpdateASCMICH ASS

During the Sim meeting today, I added parts to the ASS model so that we can also dither the BS and minimize the power at AS. 

ASS screen has been updated. 

Model changes required a new sender from LSC for ASDC, so both LSC and ASS were compiled, installed and restarted.  Also on LSC, I added AS110 I&Q to DQ channels, since we haven't been recording them in the past.

I have done a quick trial in MICH-only lock, and it seems to work.  Gain of 10 for both Pit and Yaw servos. 

  10502   Fri Sep 12 14:11:17 2014 ericqUpdateLSCDRMI locking

We need to get an interferometric estimation of the SRC length error / SRC sideband splitting, because if the 7.5cm hand-measured error is true, it looks like it might be hard to control the DRMI on 3F. 


I did some DRMI sensing simulations, to get an idea if sensing matrix elements might change as the CARM offset changes. Last night, I tried just going to zero CARM offset on ALS, and was having problems keeping the PRMI locked on REFL33, so I wanted to confirm that it should at least work in theory. 

Thus, I simulated what happens to the sensing matrix element in the vertex DoFs as the CARM offset is reduced, in both the PR and SR cases. I normalized all of the elements to PRCL at zero carm offset, to get an idea of what the good relative gains should be for MICH and SRCL. 

In the end, there don't seem to be significant DC gain changes, or demod angle fluctuations, in either the PRFPMI or DRFPMI case, as the CARM offset changes, which is good.

However, the SRC length as hand-measured, seems to mess up the MICH angle in the DRFPMI case, and really lowers the SRCL signal amplitude. 

To be fair, past efforts of simulating demodulation angles haven't always been borne out on the IFO, so we should still forge ahead experimentally until it becomes apparent that there is a real problem. 


Here are the simulations for the IFO as-is:

(A note on the plots. Though they kind of look like Bodes, they're just the sensing element represented as a complex number in the I-Q plane,I being phase=0 and Q = 90)

dcPRCL3F.pdfdcMICH3F.pdfdcSRCL3F.pdf

All three signals are along the I axis in the DRMI case, which seems like it would be tough to control, since we only have 2 3F diodes... We've been using REFL33Q when PRMIing, which is simulated at around 45 deg; it should be easy to verify this empirically. 


Here are the same plots with the SRC length corrected. Now MICH shows up mostly in the Q phase as desired in the DRMI case. SRCL in REFL165 also wins 20dB of optical gain, as well. 

dcPRCL3F_correctSRC.pdfdcMICH3F_correctSRC.pdfdcSRCL3F_correctSRC.pdf

 


To drive the point home, here's a simulated scan of AS110 and REFL55 Q to show the effect of the measured length error:

SRClengthEffects.pdf

 

  10501   Fri Sep 12 12:00:59 2014 ericqUpdateLSCDRMI locking

REFL11 I, as seen in digital land, is connected to the slow output of the CM board. I tuned the demod angle of the REFL11 demodulator board by cable length back in ELOG 9850. It would be good to check that the phase is still good. If the CM board gains are at 0dB, we should be able to used the digital angle adjustment as normal. 

  10500   Fri Sep 12 11:25:42 2014 KojiUpdateLSCDRMI locking

This is great.

And I got confused. Is REFL11 going through the CM board?
If so how the demod phase for REFL11 take an effect for the sensing?

Maybe I understood. CM SERVO SLOW has been connected to REFL11I? whitening.
Therefore using REFL11 in the CM SERVO gives us REFL11I at the usual channels.
And then how can we ensure the gain matching between I & Q?

Then is the next step 3f DRMI? How is REFL165 healthy?
I also wonder how the relative phase and modulation depths improves the sensing matrix.

  10499   Fri Sep 12 03:49:57 2014 ericqUpdateLSCSome more PRFPMI efforts

 Since DRMI didn't get fully commissioned, I tried my hand at PRFPMI locking with the newly improved ALS performance. 

ALS seemed reliable, I think my main limiting factor was the PRMI locking. We should set up a restore script for PRFPMI that is a superset of the ALS CARM DARM, because the current restore script doesn't put all the vertex settings back, so I was trying to lock for a while without the FM boosts on PRCL and MICH, which really hurt my stability. 

Transitioning to SqrtInv works fine; a couple of times I've gotten to arm power of ~10, and have been able to sit there for a while as I set up excitation line comparisons with the CM board's REFLDC, but the PRC would always lose it before I did anything interesting. 

The PRMI locks with a reasonable MICH offset, I found that adding a offset of 20 to 40 makes the AS spot visibly dimmer, and ASDC falls to ~0.05 from .1-.2. 

I looked into adding a boost to the CARM loop after transitioning to sqrtInv, but we only have 30 degrees of margin, and the error signal is already fairly white, so there isn't much to do, really. 

The ALS locking script is sporadically hanging a fair while, as well, which is strange. Otherwise, not much to report...

  10498   Fri Sep 12 00:40:23 2014 JenneUpdateLSCDRMI locking

 Tonight I worked on DRMI locking.  

I think the reason the May2014 DRMI recipe wasn't working for me is because I wasn't including the REFL11 -> SRCL element.  I had left it out because (a) I didn't think we should need it and (b) REFL11 is going through the CM board.  

Tonight, I flipped the switch on the CM screen so that OUT2 was seeing REFL11I, not REFLDC, so I had REFL11 in the usual place.  I reset the demod phase, since we had left it at zero for CM stuff. 

Setting demod phases for PRMI:

I locked PRMI on sideband, REFL 33 I&Q and drove PRM.  REFL55 was at 55deg, and I changed it to 33deg to minimize the peak in the Q-phase.  REFL11 was a 0deg, and I set it to 17deg.  I also checked the AS55 phase in the MICH-only case, and changed it from 14.75deg to 24.75 deg.

The May 2014 recipe (elog 9968) calls for adding 25 degrees to the REFL55 phase, so I put REFL55 at 58deg for DRMI locking.

After that, using the parameters in the May2014 recipe, the DRMI just locked.  Awesome!

I checked the demod phases with DRMI lock.  REFL11 stays at 17 degrees.  If I actuate the SRM, I get the largest peak in the I-phase of REFL55 with a phase of -143deg, but the acquisition is best with phase around 55deg.  [Note, as Q points out, I wonder if SRCL is mostly locked with REFL11I for some magical reason, which is why it didn't matter so much that I put a sign flip into REFL55...I wonder if fixing our macroscopic length offset in SRCL will fix this].  I also changed the REFL165 phase from -155.5deg to +145deg.

By looking at transfer functions at an excitation frequency, I expected that I should be able to hold SRCL and MICH on REFL165, with matrix elements -0.085 for REFL165I->SRCL and -0.23 for REFL165Q->MICH.  I was not able to acquire with these values, nor was I able to ramp the matrix elements while keeping lock.

So, I tried moving PRCL to REFL33I, which did work.  I used 1.245*REFL33I->PRCL, but left SRCL and MICH on REFL55 I&Q, with the REFL11I->SRCL element also there. This is where I started trying to get rid of the REFL11I element, but couldn't maintain lock most times, and could never acquire lock without it.

Next up, checking the MICH->SRCL coupling due to the output matrix.  I did as Koji did in elog 8816 , but first I copied the notches in FM10 of MICH over to PRCL and SRCL (old notch freqs were SRCL=566.1Hz, PRCL=675.1Hz, now they're all 475.1Hz).  I drove BS, and checked that the PRM element minimized the peak in REFL33I, the PRCL error signal.  I also added an SRM element to reduce the peak in REFL55I, the SRCL error signal.  I ended up with 0.5*BS, -0.284*PRM, -1.5*SRM for MICH drive, and unity in the PRM and SRM elements for PRCL and SRCL, respectively.

I measured the SRCL open loop gain, and the UGF was pretty low, so I increased the SRCL gain from 0.2 to 0.5 to make the UGF be around 70Hz.  I measured PRCL and MICH also, and they matched their references.

I worked a little bit on trying to remove REFL11 from the SRCL error signal, but didn't get anywhere.  I'm leaving the IFO to Q for the rest of the night.

To sum up, here is the set of parameters that worked for DRMI locking.  (These are saved as the template on the IFO Config screen.):

DEMOD PHASES:

REFL11:  17 deg

REFL33: 140.5 deg (not changed tonight)

REFL55: 58 deg  (58deg for DRMI, 33deg for PRMI)

REFL165: 145 deg

AS55: 24.75 deg

INPUT MATRIX

MICH = 0.15 * REFL55Q

PRCL = 1.245 * REFL33I

SRCL = -0.09 * REFL11I + 1.0 * REFL55I

DOF Triggers

MICH, PRCL, SRCL all on POP22I, 50:10

GAINS

MICH = 1.0

PRCL = -0.02

SRCL = 0.5

FM triggers 

MICH:  35:2, 2 sec delay, FM 2, 3, 6, 9

PRCL:  35:2, 0.5 sec delay, FM 2, 3, 6, 9

SRCL:  35:2, 5 sec delay, FM 3, 6, 9  (always lose lock trying to engage FM2).

OUTPUT MATRIX

MICH = 0.5 * BS +  (-0.284)*PRM + (-1.5)*SRM

PRCL = 1*PRM

SRCL = 1*SRM

  10497   Fri Sep 12 00:28:04 2014 ericqUpdateLSCHoly sensitivity, Batman!

I took a quick measurement of the ALS stability, using POX and POY as out of loop sensors, using a CARM calibration line to line POX and POY up to the calibrated PHASE_OUT channels at 503Hz. 

  • X arm RMS ~1kHz
    • Could use more low frequency suppression
  • Y arm RMS ~200Hz

ALSoutOfLoop.pdf

  10496   Thu Sep 11 17:12:42 2014 SteveUpdateSUSOpLev test: old SP qpd connected

IP POS cable was swapped with old SP-QPD sn222 at the LSC rack.  So there is NO IP POS temporarily.

This QPDsn222 will be used the HeNe oplev test for aLIGO

 

  10495   Thu Sep 11 15:50:49 2014 SteveUpdateSUSETMX damping restored
  10494   Thu Sep 11 02:08:32 2014 JenneUpdateLSCHigher transmission powers

No breakthroughs tonight. 

DRMI didn't want to lock with either the recipe that we used a year ago (elog 9116) or that was used in May (elog 9968).  Being lazy and sleepy, I chickened out and went back to PRFPMI locking. 

Many attempts, I'll highlight 2 here.

(1) I had done the CARM -> sqrtInvTrans transition, and reduced the CARM offset to arm powers of about 7, and lost lock.  I don't remember now if I was trying to transition DARM to AS55, or if I was just prepping (measuring error signal ratio and relative sign).

Zoom_TRXTRYat7_CARMsqrtInvTrans_DARMalsdiff.png

(2) I stopped the carm_cm_up script just before it wanted to do the CARM -> sqrtInvTrans transition, and stayed with CARM and DARM both on ALS.  I got to reasonably high powers, and was measuring the error signal ratios I needed for CARM -> REFL DC and DARM -> AS55.  Things were too noisy to get good coherence for the DARM coefficient, but I thought I was in good shape to transition CARM to REFL DC (which looks like REFL11I, since REFLDC goes to the CM board, and the OUT2 of that board is used to monitor the input to the board. )  Anyhow, I set the offset such that it matched my current CARM offset value, and started the transition, but lost lock about halfway through.  CARM started ringing up here, and I think that's what caused this lockloss.  Could have been the CARM peak, which I wasn't considering / remembering at the time.

ALSonly_TryingTransitionStraightToREFLDC.png

Daytime activity for Thurs:  Lock DRMI, maybe first on 1f signals, but then also on 3f signals.

  10493   Thu Sep 11 00:27:39 2014 JenneUpdateSUSSRM oplev touch-up

The sum vs. pitch and yaw signals for the SRM QPD weren't making sense to me - centering on the PD lowered the sum, etc.  So, I had a look at the SRM oplev setup. 

The beam going in to the chamber looked fine, but the beam coming out was weird, like it was being clipped, or diffracted off of a sharp edge.  The beam was spread out in yaw over almost 1cm as seen by eye.  I looked into the vacuum window, and the beam was sitting on the edge of one of the in-vac steering optics.  So, I adjusted the yaw of the beam-launching optic on the out of vac table so that I was roughly centered on both of the in-vac SRM steering mirrors.  This required moving the first out of vac mirror for the SRM oplev path on the way to the QPD to move a small amount to one side, since the beam was near-ish the edge of the optic.  I then centered the beam on the oplev (I had the SRM roughly aligned already). 

Now the SRM oplev makes more sense to me.  I have turned on FMs 1, 2, 5, 9 to match ITMY's loop shape.  I have set the gains to -10 for pitch and +10 for yaw, to make the upper UGF about 6 Hz.

  10492   Wed Sep 10 22:17:29 2014 KojiSummaryLSCX/Y green beat mode overlap measurement

[Koji Manasa]

We made quantitative inspection of the X/Y green beat setup on the PSL table.

DC output of the BBPD for each arm was measured by blockiing the beams at either or both side of the recombination BS.

The power over lap for the X arm beat note setup was 7.8% and is now 53%.
There is 3dB of headroom for the improvement of the mode overlap.

The power over lap for the Y arm beat note setup was 1.2% and is now 35%.
There is 4dB of headroom for the improvement of the mode overlap.

The RF analyzer monitor for the beat power is about 10dB lower than expected. Can we explain this only by the cable loss?
If not it there something causing the big attenuation?


             XARM   YARM
o BBPD DC output (mV)

 V_DARK:   -  3.3  + 1.9
 V_PSL:    +  4.3  +22.5
 V_ARM:    +187.0  + 8.4


o BBPD DC photocurrent (uA)

I_DC = V_DC / R_DC ... R_DC: DC transimpedance (2kOhm)

 I_PSL:       3.8   10.3
 I_ARM:      95.0    3.3


o Expected beat note amplitude
I_beat_full = I1 + I2 + 2 sqrt(e I1 I2) cos(w t) ... e: mode overwrap (in power)

I_beat_RF = 2 sqrt(e I1 I2)

V_RF = 2 R sqrt(e I1 I2) ... R: RF transimpedance (2kOhm)

P_RF = V_RF^2/2/50 [Watt]
     = 10 log10(V_RF^2/2/50*1000) [dBm]

     = 10 log10(e I1 I2) + 82.0412 [dBm]
     = 10 log10(e) +10 log10(I1 I2) + 82.0412 [dBm]


for e=1, the expected RF power at the PDs [dBm]
 P_RF:      -12.4  -22.6


o Measured beat note power (before the alignment)     
 P_RF:      -23.5  -41.7  [dBm] (38.3MHz and 34.4MHz) 
    e:        7.8    1.2  [%]                         
o Measured beat note power (after the alignment)      
 P_RF:      -15.2  -27.1  [dBm] (26.6MHz and 26.8MHz) 
    e:       53     35    [%]                         

Measured beat note power at the RF analyzer in the control room
 P_CR:      -25    -20    [dBm]
Expected    -17    - 9    [dBm]

Expected Power:
Pin + External Amp Gain (0dB for X, 20dB for Y)
    - Isolation trans (1dB)
    + GAV81 amp (10dB)
    - Coupler (10.5dB)


  10491   Wed Sep 10 21:05:43 2014 JenneUpdateLSCHoly sensitivity, Batman!

Koji and Manasa did some work on the PSL green situation today (Koji is still writing that log post up), but I just measured the Yarm out of loop sensitivity, and WOAH. 

The beat is -11.5dBm at 42.8 MHz.  Koji said the sweet spot is around 30 MHz.  The out of loop sensitivity is 400 Hz RMS!  Something to note is that the Y beatnote still has a 20dB amplifier before going to the beatbox, but the X does not.  We had been worried about saturation issues with the X, so we took out the amplifier.  However, I might put it back if we win big like this.

Recall from elog 10462 that I had saved a reference of the out of loop noise for both X and Y, but Y was much noisier than X.  The references below are from that elog, and the new Y is in dark blue. (Edit, 9:18pm, updated plot measuring down to 0.01Hz.  This is the new reference on the ALS_outOfLoop_Ref.xml template).

 ALS_Y_outOfLoop_10Sept2014.pdf

EDIT:  (Don't worry, I'm going to measure X too, but right now the beam overlap on the camera is not good, as if something drifted after Koji and Manasa closed up the PSL table)

Touched up the alignment for X on the PSL table.  Current beatnotes are:  [Y, -13.5 dBm, 74.1 MHz], [X, -22 dBm, 13.9 MHz].  Red is the current X out of loop, and I've saved it as the new X reference on the template.

ALS_XY_outOfLoop_10Sept2014.pdf

  10490   Wed Sep 10 20:24:00 2014 JenneUpdateElectronicsPOY RF cable loose

Sitting down to work on the IFO, I couldn't lock the Yarm.  I looked at the error signal as well as the transmission on Dataviewer, as usual, and saw that the POY error signal was almost non-existant. 

Since there was work on the POY table today (Steve removed the oplev test setup, elog 10489 and Q centered the SRM oplev after doing SRMI alignment, no elog yet), I went out to have a look at the table. 

There was nothing occluding the POY beam, which I traced back to the edge of the table.  The beam looked nice and round, so I decided that wasn't it.  I jiggled the PD cables, and lo and behold, the POY RF out cable almost came off in my hand it was so loose.  My suspicion is that whomever was the last to put the POY RF out back didn't tighten the cable and then the work today jiggled the cable loose.  I tightened the cable, and by the time I was back to the control room the arm was locked and Koji was already running the alignment scripts.

  10489   Wed Sep 10 15:31:16 2014 SteveUpdateSUSHeNe laser test preparation

Quote:

Quote:

 

 SRM as set up in Atm4  26,000 count compared with ETMY oplev servo in operation 7,500 counts for 3 days

Next steps: measure beam size at qpd,

                  place qpd on translation stage for calibration,

                  change 1103P mount to single one

 SRM qpd is installed on translation stage and the shims removed from laser V mounts.

The ETMY oplev servo is on.

SRM oplev servo:  100 microrad/count is an estimate, not calibrated one.

 SRM qpd is back to its normal position. The mount base is still on delrin base. SRM and ITMY need centering.

Tomorrow I will set up the HeNe laser test at the SP table with Ontrack qpd

ETMY oplev servo on. SRM qpd with ND1 ------no component------- 1103P

Attachment 1: SRMqpdisBack0nSRMsus.png
SRMqpdisBack0nSRMsus.png
Attachment 2: slowVSfast.png
slowVSfast.png
  10488   Wed Sep 10 14:58:58 2014 JenneUpdateSUSOpLev test: New channels

Steve and EricG are moving their oplev test for aLIGO over to the SP table, so that we can have the SRM optical lever back.  

I have pulled out an Ontrak PSM2-10 position sensor and accompanying driver for the sensor.  This, like the POP QPD, has BNC outputs that we can take straight to the ADC.

In the c1pem model I have created 3 new filter modules:  C1:PEM-OLTEST_X, C1:PEM-OLTEST_Y, and C1:PEM-OLTEST_SUM.  I built, installed and restarted the model, and also restarted the daqd process on the frame builder.  On the AA breakout board on the 1X7 rack, these correspond to:

BNC # 29 = OLTEST_X

BNC # 30 = OLTEST_Y

BNC # 31 = OLTEST_SUM

By putting 1Vpp, 0.1Hz into each of these channels one at a time, I see on StripTool that they correspond as I expect.

Everything should be plug-and-play at this point, as soon as Steve is ready with the hardware.

  10487   Wed Sep 10 10:49:39 2014 ManasaUpdateLSCY arm green + PSL green mode overlap

Quote:

Manasa, can you please estimate what kind of mode matching we have on the PSL table between the arm greens and the PSL green?  We *do not* want to touch any optics at this point.  Just stick in a power meter to see how much power we're getting from each beam, and then think about the peak height we see, and what that might tell us about our mode overlap.  If we determine it is total crap, we can think about measuring the beams that go either toward the camera, or the DC PDs, since neither of those paths require careful alignment, and they are already picked off from the main beatnote path.  But first, what is our current efficiency?  Yarm is first, then Xarm, since Yarm seems worse (peak height is larger for non-00 modes!)

Estimate loss along the Y arm beat path:

1. Measured the beam powers (before the beam combiner): 
Y Arm green = 35 uW
Y PSL green = 90 uW

==> Pbeat ~ 2 * sqrt (35 uW * 90 uW) ~ 112 uW

2. Expected power of RF signal
Assuming the PD to have transimpedance ~ 2kV/A and responsivity ~ 0.3A/W, 
the expected power of the RF signal = (Pbeat * Transimpledance* Responsivity)^2 / (2 * 50ohm) ~ 45uW = -13.5 dBm

3. Measured power of Y arm beat signal

Turned OFF the beat PDs and rerouted the RF cables such that the spectrum analyzer was reading the RF signal from the Y beat PD itself (without any amplifiers or the beat box itself in the path).  
Turned ON the beat PDs and the Y arm beat signal power on the spectrum analyzer measured -58dBm
Even if we consider for losses along the length of the cables, we are still at a very bad state. 

4. Bad mode matching??
I don't think mode matching is our main problem here.
Toggling the shutter several times, even with the non-00 modes, the maximum beat power we can see is -50dBm which is still very far from the actual expected value.

  10486   Wed Sep 10 02:59:42 2014 ericqUpdateLSCLocking activities - nothing new :(

Some small things I did tonight which did little to nothing to help:

  • I reset the offsets in the SQRTINV FMs to try and match the DC level of the ALS CARM error signal as best as possible, to avoid moving away from the set-point too much, as I was worried we were wandering into regions of too low optical gain. 
  • I turned off the WFS, and hand tweaked the MC alignment. The WFS loops / matrices definitely have some room for improvement, and I was worried that excess angular motion of the MC was coupling into CARM. MC refl is much calmer in the last ~1.5 hrs since I turned off the WFS. 

My main concern with tonights situation was the huge low frequency fluctuations of TRY while CARM/DARM locked on ALS. We saw this being very smooth very recently, but when one arm is fluctuating by multiple line widths, it isn't surprising that locks aren't stable. I want to know why the out of loop stability is so unpredictable. 

  10485   Wed Sep 10 02:53:32 2014 JenneUpdateLSCLocking activities - nothing new :(

[Jenne, EricQ]

No major progress today.  

I fixed a bug in my lockloss script that was asking it to start gathering data just after the lockloss, rather than some seconds beforehand.  Ooops.  Anyhow, with this handy-dandy plotting, I still don't know why we are losing lock when we have PRMI on REFL33, CARM on sqrtInvTrans, and DARM on AS55.  I don't see any oscillations, just the arm power drops off, and a moment later the POP power drops.  

For example, here is one of the best states we got to tonight.  Data for this is in ..../scripts/LSC/LocklossData/1094369700 .  You can re-create the plot by going to ..../scripts/LSC/LocklossData/  and doing ./PlotLockloss.py 1094369700 .  We had set the triggers for the trans PD/QPD such that we were using the QPD transmission signals the whole time (above trans of 0.2).  We saw that the noise at high frequency during low transmission powers for sqrtInvTrans as an error signal was higher using the QPDs than with the Thorlabs PDs, but that both cases are below the noise for ALS.  The arm powers were pretty steady above 3 for the last bit of this lock stretch.  I lost lock while trying to transition DARM over to AS55Q.  CARM was on sqrtInvTrans(QPDs), PRMI on REFL33 I&Q as usual.

LocklossZoom.png

 


Other things from this evening:

* When I was starting, I saw that when I locked the PRMI, the PRM was oscillating in pitch. Oscillation only happened when PRM pitch oplev was on.  I'm not sure what could have changed to make the oplev loop unstable, but the gain was 7.0, and now I have left it at 5.0.

* I recentered the PRM and ITMY oplevs.

* Plugged in the Yend PDH error monitor and pzt output monitors, since I forgot them last week.  Hopefully this will allow the Yend SLOW servo to work, and keep us away from the limits of the PZT range.

  10484   Wed Sep 10 02:52:05 2014 ericqUpdateCDSOTTAVIA lost network connection

I checked chiara's tables, all seemed fine. I switched ethernet cables from the black one labelled "allegra," which seemed maybe fragile, for the teal one that may have been chiara's old ethernet cable. It's back on the network now; hopefully it lasts. 

  10483   Wed Sep 10 02:35:55 2014 ranaUpdateCDSOTTAVIA lost network connection

Quote:

Today the network connection of OTTAVIA was sporadic.

Then in the evening OTTAVIA lost completely it. I tried jiggle the cables to recover it, but in vain.

We wonder if the network card (on-board one) has an issue.

 I would also suspect IP conflicts; I had temporarily put the iMac on the Ottavia IP wire a few weeks ago. Hopefully its not back on there.

  10482   Wed Sep 10 02:35:54 2014 JenneHowToTreasureSecret scripts, revealed!

 I hereby confess to having a secret script.  But it is secret no longer!

It's a "goLock" script, and it is now in the path from any terminal.  It kills any open medm sessions (to clean up desktops), and then opens a palette of screens that I find useful.  It also starts up the CARM and DARM ALS watch scripts, and the toggle shutter scripts.  It then leaves the terminal in .../scripts/PRFPMI/ , which is where the carm_cm_up.sh script that we've been using lives.

I also made tonight a "goHome" script, but all that one does so far is set the LSC mode to OFF.  The other thing that this could / should do is restore all optics so we don't have hysteresis problems.

Also, also, my "new" misalign / restore scripts had a bug, in that they were always switching oplevs for the PRM, no matter what optic was requested.  This sometimes caused the PRM oplev to be engaged while the optic was misaligned, so the PRM would get rung up.  This has been fixed.

ELOG V3.1.3-